# Revisiting Few-Shot Object Detection with Vision-Language Models

### v2

@@ Appendix D RegionCLIP Experiments
Detic (Zero-Shot) [49] 14.26 27.28 16.88 2.36
GroundingDINO (Zero-Shot) [23] 11.44 17.42 16.13 3.38
RegionCLIP (Zero-Shot) w/ LVIS-RPN [47] 2.34 3.33 3.87 0.22
RegionCLIP (Zero-Shot) w/ Detic-RPN [47] 3.79 6.68 3.91 1.12
RegionCLIP (Zero-Shot) w/ Detic-RPN, 0.5 [47] 7.64 12.81 9.57 1.88
RegionCLIP (Zero-Shot) w/ GT-RPN [47] 26.44 45.33 31.83 3.92
Appendix E NuImages Per-Split Breakdown
We provide per-split breakdowns for the nuImage experiments in Tables 11, 12, and 13.
Table 11: 5-shot nuImages Per-Split Performance
Approach Split 1 Split 2 Split 3 Split 4 Split 5
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-tuning 15.07 24.89 18.82 4.58 13.27 20.83 17.54 3.58 15.15 26.55 19.23 2.93 14.82 25.47 19.10 3.20 14.76 23.96 19.45 3.69
FedLoss 15.27 26.59 19.25 3.41 15.08 27.23 18.59 2.80 15.22 26.69 19.34 2.97 14.72 25.42 19.44 2.65 15.03 25.87 19.39 3.01
InvFedLoss 15.36 28.54 18.59 2.77 15.07 28.17 18.04 2.71 15.40 27.91 18.96 2.94 15.04 26.97 18.95 2.80 15.55 28.28 18.93 3.12
Pseudo-Negatives 15.87 28.59 19.06 3.67 15.88 28.70 19.28 3.30 16.12 27.94 20.01 3.84 15.76 28.00 19.65 3.37 16.21 28.43 19.90 3.80
True Negatives (Oracle) 16.27 28.55 19.35 4.56 15.92 29.03 19.24 3.22 15.97 28.81 19.42 3.29 15.92 28.61 19.57 3.41 16.44 28.91 20.1 3.92
Exh. Annotations (Oracle) 17.59 32.63 20.3 3.97 17.71 32.91 21.23 3.31 17.13 32.03 20.42 3.02 17.81 32.86 21.51 3.41 17.72 32.88 20.99 3.48
Approach Split 6 Split 7 Split 8 Split 9 Split 10
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-Tuning 14.94 24.77 19.37 3.84 14.58 25.02 18.26 3.54 15.15 26.56 19.14 3.38 15.39 26.42 18.90 4.20 14.58 27.13 16.73 3.38
FedLoss 15.27 26.30 19.82 3.06 14.88 27.23 18.41 2.59 14.85 25.27 19.68 3.04 15.25 26.40 19.15 3.58 14.79 27.09 17.61 3.16
InvFedLoss 14.65 23.63 19.56 3.67 14.21 23.34 18.73 3.55 14.86 24.71 19.82 3.44 15.74 27.73 19.13 3.97 15.29 28.43 18.19 3.02
Pseudo-Negatives 16.30 27.66 20.89 3.82 15.82 28.69 19.00 3.47 15.82 27.93 20.11 3.28 15.73 29.08 18.95 3.04 15.87 29.67 18.46 3.37
True Negatives (Oracle) 16.46 29.5 20.38 3.43 15.95 29.54 18.94 3.26 16.23 28.4 20.14 4.18 16.52 29.4 19.77 4.12 15.8 29.3 18.31 3.6
Exh. Annotations (Oracle) 18.07 32.93 21.84 3.7 17.41 32.67 20.71 3.17 17.54 32.81 20.77 3.39 17.4 32.45 20.26 3.69 17.54 32.46 21.27 3.14
Table 12: 10-shot nuImages Per-Split Performance
Approach Split 1 Split 2 Split 3 Split 4 Split 5
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-Tuning 15.77 26.63 20.13 3.87 15.31 26.85 19.46 3.03 15.64 27.32 20.08 2.91 15.78 25.16 20.79 4.44 15.93 27.34 20.28 3.60
FedLoss 15.66 27.74 20.05 2.83 15.53 25.71 20.84 3.07 15.62 28.49 19.69 2.43 15.63 27.14 20.31 2.91 15.72 25.22 21.37 3.54
InvFedLoss 15.96 27.22 20.51 3.66 15.80 26.05 20.88 3.54 15.85 28.71 19.97 2.62 16.04 27.43 20.49 3.68 16.26 27.05 21.22 3.80
Pseudo-Negatives 16.78 29.33 20.64 4.14 16.58 29.73 20.53 3.36 16.51 28.95 20.72 3.44 16.95 28.73 21.49 4.12 17.08 29.84 21.22 3.90
True Negatives (Oracle) 16.91 29.67 20.67 4.19 16.92 29.75 21.08 3.73 16.57 29.69 20.38 3.4 17.36 29.96 21.24 4.67 17.52 29.46 22.13 4.61
Exh. Annotations (Oracle) 18.03 33.17 21.33 3.77 18.68 33.98 22.52 3.88 18.3 33.2 22.37 3.45 18.44 33.47 22 4 19.08 33.71 23.36 4.42
Approach Split 6 Split 7 Split 8 Split 9 Split 10
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-Tuning 15.22 23.86 21.10 3.65 15.00 22.86 19.67 5.07 15.67 27.42 19.67 3.55 15.25 25.67 19.61 3.60 15.73 27.02 18.52 5.10
FedLoss 15.79 26.93 21.08 2.82 15.57 27.22 19.76 3.17 15.47 27.98 19.46 2.73 15.32 27.02 19.92 2.44 15.40 28.60 18.45 2.92
InvFedLoss 16.07 27.49 21.09 3.15 15.93 28.80 19.43 3.31 15.69 28.38 19.45 3.04 15.29 25.94 20.12 2.94 15.99 28.50 18.72 4.43
Pseudo-Negatives 16.60 29.51 20.83 3.33 16.50 27.91 20.24 4.82 16.51 29.26 20.61 3.50 16.37 28.86 20.75 3.09 16.82 29.37 19.56 5.27
True Negatives (Oracle) 17.06 29.29 22.12 3.59 16.43 29.9 19.55 3.76 17.04 29.85 20.77 4.46 17.11 29.38 21.48 4.02 16.96 29.04 19.86 5.63
Exh. Annotations (Oracle) 18.62 33.65 23.09 3.45 18.75 33.83 22.2 4.51 17.99 33.11 21.45 3.78 18.33 33.18 22.47 3.51 18.94 33.77 22.67 4.55
Table 13: 30-shot nuImages Per-Split Performance
Approach Split 1 Split 2 Split 3 Split 4 Split 5
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-Tuning 16.71 28.14 21.37 4.14 15.94 26.49 21.04 3.67 16.95 27.73 22.11 4.24 16.33 28.71 20.41 3.52 17.46 28.50 22.64 4.68
FedLoss 16.20 28.86 20.76 2.75 15.48 27.30 20.41 2.31 16.59 28.80 21.71 2.82 16.62 28.85 21.45 3.09 17.15 29.59 21.94 3.53
InvFedLoss 16.47 29.26 20.70 3.28 16.09 27.10 21.52 3.09 16.78 29.52 21.18 3.40 16.97 29.25 21.39 3.90 17.57 29.84 22.14 4.35
Pseudo-Negatives 18.14 30.45 22.33 5.38 17.34 29.24 22.53 3.91 17.90 31.09 22.03 4.42 17.83 30.44 22.34 4.35 18.33 30.78 23.04 4.87
True Negatives (Oracle) 18.34 30.55 22.52 5.68 17.82 29.46 22.89 4.83 17.76 31.09 21.93 4.15 18.36 30.99 22.61 5.18 18.67 31.03 23.59 5.18
Exh. Annotations (Oracle) 19.48 34.22 23.24 5.12 19.43 34.39 23.81 4.47 19.33 34.3 23.97 3.98 19.21 34.18 23.44 4.15 19.6 34.2 24.22 4.62
Approach Split 6 Split 7 Split 8 Split 9 Split 10
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-Tuning 16.66 28.51 21.51 3.59 17.30 28.02 21.89 5.53 17.19 27.68 22.42 5.01 17.05 28.11 21.84 4.56 16.68 27.12 20.67 5.57
FedLoss 17.01 28.95 22.57 3.08 16.22 29.48 19.98 3.14 16.46 29.15 21.14 3.03 16.47 29.46 20.89 2.84 16.30 28.33 20.56 3.59
InvFedLoss 17.03 29.64 22.03 3.16 17.36 29.69 21.66 4.60 17.05 29.18 21.90 4.02 16.65 29.73 20.88 3.18 16.82 28.65 20.66 4.77
Pseudo-Negatives 18.08 30.58 23.00 4.39 18.05 30.20 22.37 5.43 17.61 30.47 22.01 4.33 17.87 30.13 22.50 4.64 17.56 29.77 21.34 5.29
True Negatives (Oracle) 17.79 30.57 22.92 3.71 18.44 30.47 22.66 6.07 18.3 30.83 22.72 5.34 18.39 30.74 23.02 5.08 18.07 30.35 21.8 5.88
Exh. Annotations (Oracle) 19.32 34.21 24.33 3.71 19.47 34.47 23.38 4.91 19.4 34.17 23.7 4.66 19.47 34.12 23.94 4.57 19.34 34.17 23.4 4.65

### v3

@@ Appendix E Impact of Box-Level Supervision for Foundational FSOD
Detic (Zero-Shot) [zhou2022detecting] 14.26 27.28 15.15 2.36
GroundingDINO (Zero-Shot) [liu2023grounding] 11.44 17.42 14.08 3.38
RegionCLIP (Zero-Shot) w/ LVIS-RPN [zhong2022regionclip] 2.34 3.33 3.45 0.22
RegionCLIP (Zero-Shot) w/ Detic-RPN [zhong2022regionclip] 3.79 6.68 4.01 1.12
RegionCLIP (Zero-Shot) w/ Detic-RPN, 0.5 [zhong2022regionclip] 7.64 12.81 8.88 1.88
RegionCLIP (Zero-Shot) w/ GT-RPN [zhong2022regionclip] 26.44 45.33 32.25 3.92
Appendix F NuImages Annotator Instructions
We present an example of the nuImages annotator instructions below. Notably, such annotator instructons are naturally few-shot (e.g. providing a few visual and textual examples describing the target concept), multi-modal, and contain both positive and negative examples. Our proposed Foundational FSOD benchmark, and pseudo-negative federated loss facilitate future work in leveraging rich annotator descriptions, allowing us to “align” VLMs much like how annotators must be “aligned” to subtle aspects of the target class.
Figure 6: NuImages Annotator Instructions. We include the multi-modal annotator instructions barrier. Our proposed setup allows FSOD methods to learn such multi-modal examples, similar to how human annotators are taught the labeling policy. Importantly, annotators can also be provided with negative examples (in red) for classes, i.e what NOT to label for a certain class. Crucially, our proposed fine-tuning with pseudo-negatives can easily accommodate such negative examples within the proposed setup.
Table 10: Empirical Analysis of Baselines (5-Shots) on nuImages.
Approach Backbone Pre-Train Average Precision (AP)
Data All Many Med Few
Zero-Shot Detection
   RegionCLIP [zhong2022regionclip] RN50 CC3M 2.50 3.20 3.80 0.40
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 14.40 25.83 16.59 2.32
   GroundingDINO [liu2023grounding] SWIN-T Objects365,GoldG,Cap4M 12.05 17.29 15.45 3.72
   GLIP [li2021grounded] SWIN-L FourODs,GoldG,Cap24M 17.01 23.36 19.86 8.40
   MQ-GLIP-Text [xu2024multi] SWIN-L Objects365,FourODs,GoldG,Cap24M 17.01 23.36 19.85 8.41
Prompt Engineering
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 14.92 26.48 17.29 2.53
   GLIP [li2021grounded] SWIN-L FourODs,GoldG,Cap24M 17.15 23.82 19.36 9.02
Standard Fine-Tuning
   RegionCLIP [zhong2022regionclip] RN50 CC3M 3.84 6.13 5.07 0.49
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 15.12 22.74 18.99 4.25
Federated Fine-Tuning (Ours)
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 16.58 27.12 19.71 4.13
   Detic [zhou2022detecting] w/ Prompt Engineering SWIN-B LVIS, COCO, IN-21K 16.96 27.89 19.94 4.37
Language Prompt Tuning
   GLIP [li2021grounded] SWIN-L FourODs,GoldG,Cap24M 17.79 21.07 22.87 9.12
Visual Prompting
   MQ-GLIP-Image [xu2024multi] SWIN-L Objects365,FourODs,GoldG,Cap24M 13.42 23.05 15.00 3.54
Multi-Modal Prompting
   MQ-GLIP [xu2024multi] SWIN-L Objects365,FourODs,GoldG,Cap24M 21.45 32.23 23.31 10.30
Multi-Modal Chat Assistants
   GPT-4o Zero-Shot Classification [achiam2023gpt] Private Private 9.95 16.81 12.11 1.71
Table 11: Empirical Analysis of Baselines (30-Shots) on nuImages.
Approach Backbone Pre-Train Average Precision (AP)
Data All Many Med Few
Zero-Shot Detection
   RegionCLIP [zhong2022regionclip] RN50 CC3M 2.50 3.20 3.80 0.40
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 14.40 25.83 16.59 2.32
   GroundingDINO [liu2023grounding] SWIN-T Objects365,GoldG,Cap4M 12.05 17.29 15.45 3.72
   GLIP [li2021grounded] SWIN-L FourODs,GoldG,Cap24M 17.01 23.36 19.86 8.40
   MQ-GLIP-Text [xu2024multi] SWIN-L Objects365,FourODs,GoldG,Cap24M 17.01 23.36 19.85 8.41
Prompt Engineering
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 14.92 26.48 17.29 2.53
   GLIP [li2021grounded] SWIN-L FourODs,GoldG,Cap24M 17.15 23.82 19.36 9.02
Standard Fine-Tuning
   RegionCLIP [zhong2022regionclip] RN50 CC3M 3.87 6.05 5.14 0.57
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 17.22 25.98 21.64 4.78
Federated Fine-Tuning (Ours)
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 18.64 29.13 22.44 5.46
   Detic [zhou2022detecting] w/ Prompt Engineering SWIN-B LVIS, COCO, IN-21K 18.67 29.13 22.43 5.57
Language Prompt Tuning
   GLIP [li2021grounded] SWIN-L FourODs,GoldG,Cap24M 20.73 24.95 25.60 11.54
Visual Prompting
   MQ-GLIP-Image [xu2024multi] SWIN-L Objects365,FourODs,GoldG,Cap24M 14.26 24.55 16.73 2.79
Multi-Modal Prompting
   MQ-GLIP [xu2024multi] SWIN-L Objects365,FourODs,GoldG,Cap24M 21.40 32.08 23.31 10.27
Multi-Modal Chat Assistants
   GPT-4o Zero-Shot Classification [achiam2023gpt] Private Private 9.95 16.81 12.11 1.71
Appendix G Empirical Analysis of Baselines (5-Shots and 30-Shots)
We evaluate all baselines for the nuImages experiments with 5-shots and 30-shots in Tables  10 and 3 respectively. We find that trends from the main paper hold. Notably, MQ-GLIP with-multi-modal prompting performs the best. However, we find that adding more examples (e.g. MQ-GLIP 5-shot vs. MQ-GLIP 30-shot) doesn’t seem to help in-context learning based methods nearly as much as gradient-based fine-tuning approaches.
Appendix H Foundational FSOD with LVIS
Table 12: LVIS Foundational FSOD Performance. We present fine-tuning results for different variants of Detic on the LVIS 10-shot dataset. We follow the standard FSOD setup and pre-train Detic on LVIS-base for fair comparison with prior work. Detic pre-trained only on LVIS-base outperforms specialized methods like TFA and DiGeo by $\sim$6 AP, without fine-tuning on rare classes. Since we keep the model backbone (ResNet-50) and pre-training data same for all methods, these performance improvements can be attributed to Detic’s CLIP-based classifier. This demonstrates that concept leakage through language significantly improve FSOD, and leveraging language cues should be embraced in data constrained settings. Naively fine-tuning Detic yields a performance drop of $AP_{f}$ and $AP_{c}$ because treating common classes as negatives in rare category federated datasets hurts performance. Instead, we find that embracing the federated nature of FSOD datasets provides consistent improvements in fine-tuning (30.0 vs. 30.8 for ResNet-50). Further, pseudo-labeling negatives in each image provides a modest improvement (30.8 vs. 31.6 for ResNet-50). Similar trends hold for the Swin backbone.
Approach 10-shots
$AP$ $AP_{f}$ $AP_{c}$ $AP_{r}$
ResNet-50 Backbone
TFA w/ fc [wang2020frustratingly] 24.1 27.9 23.9 14.9
TFA w/ cos [wang2020frustratingly] 24.4 27.7 24.3 16.9
DiGeo [ma2023digeo] 24.9 28.5 24.6 17.3
Detic (Base Only) [zhou2022detecting] 30.0 34.4 30.8 16.3
   + Fine-Tuning (Base + Novel) 30.0 33.2 31.9 15.5
   w/ FedLoss 30.8 33.9 32.7 17.4
   w/ Pseudo-Negatives 31.6 34.8 32.8 19.8
Swin Backbone
Detic (Base Only, SWIN-B) [zhou2022detecting] 35.2 38.7 36.8 21.4
   + Fine-Tuning (Base + Novel) 35.9 37.1 37.8 26.7
   w/ FedLoss 36.5 36.7 38.3 30.4
   w/ Pseudo-Negatives 37.2 37.7 38.2 32.6
MQ-GLIP-Text (SWIN-L) 35.8 40.2 33.1 33.0
MQ-GLIP-Image (SWIN-L) 28.8 33.0 26.6 25.1
MQ-GLIP (SWIN-L) 43.4 46.4 41.8 40.1
Although we use nuImages for Foundational FSOD for benchmarking in the main paper and in our competition, other datasets can still be evaluated under this framework. We include benchmarking results for LVIS below. LVIS [gupta2019lvis] re-annotates COCO images using 1,230 fine-grained classes, which are divided into frequent, common and rare based on the cardinality of each class. Frequent and common classes are combined to form LVIS-base and is used for pre-training. Rare classes are used for LVIS-novel. Following [wang2020frustratingly, ma2023digeo], we benchmark with LVIS v0.5 on publicly released data splits and report performance averaged across 3 splits for frequent, common, and rare groups ($AP_{f},AP_{c},AP_{r}$) on the LVIS val-set.
As shown in Table 12, Detic outperforms all recent FSOD baselines including DiGeo [ma2023digeo] by about $\sim$6 points on $AP_{c}$ and $AP_{f}$ and achieves $16.3$ $AP_{r}$ without ever seeing any rare class data (e.g by prompting Detic (Base Only) with the rare class names). Importantly, these performance improvements can be attributed to Detic’s CLIP-based classifier, which uses CLIP text embeddings corresponding to class names. Such embeddings are a result of large-scale pre-training, which we can effectively leverage for the few-shot task. This highlights the role of language in data-constrained settings.
Further, fine-tuning Detic with pseudo-negatives improves overall performance by $1.6$ AP ($30.0$ vs $31.6$) over naive fine-tuning. To contextualize the improvement in performance, we note that between TFA (ICML 2020) and DiGeo (CVPR 2023), the community improved on LVIS FSOD by only $0.5$ AP (cf. Table 12). Finally, we note that simply replacing the ResNet-50 backbone with a Swin-B transformer yields a sizeable $12.8$ AP improvement for rare classes ($19.8$ vs. $32.6$).
We present fine-tuning results for different variants of Detic on the LVIS 10-shot dataset. Following the standard FSOD protocol, we pre-train Detic on LVIS-base (e.g. frequent and common classes) and fine-tune on 10-shots from each class in LVIS-base and LVIS-novel. Importantly, this means that only results for $AP_{r}$ are indicative of true few-shot performance. First, we find that naively fine-tuning Detic on Base + Novel yields lower performance for $AP_{f}$ and $AP_{r}$. Intuitively, this suggests that ignoring the federated nature of FSOD datasets (e.g. by following the standard practice of assuming common classes are negatives for rare class federated datasets) hurts common class performance (cf. Table 12). Importantly, simply training with FedLoss significantly improves over naive fine-tuning, increasing $AP_{r}$ by 1.9% (15.5 vs. 17.4) and 3.7% (26.7 vs. 30.4) for the ResNet-50 and Swin backbones respectively. Further, leveraging our proposed negative pseudo-labeling strategy provides further improvements over the naive federated loss, increasing $AP_{r}$ by another 2.4% (17.4 vs. 19.8) and 3.7% (30.4 vs. 32.6) for the ResNet-50 and Swin backbones respectively. Similar to nuImages, we find that multi-modal prompting with MQ-GLIP performs the best of all baselines tested, significantly improving over MQ-GLIP-Text and MQ-GLIP-Image. We attribute MQ-GLIP’s strong performance to its bigger backbone and significantly larger pre-training dataset.
LVIS v0.5 Detic Experiment Details. We select Detic with a Resnet-50 backbone for fair comparison with prior work. We pre-train Detic on LVIS-base for $90k$ iterations with a batch size of $32$ using an AdamW optimizer and a learning rate of $32e-$ 3. All images are resized to $640\times 640$ and we also enable Repeat Factor Sampling [gupta2019lvis]. Following [wang2020frustratingly], we sample up to $10$ shots for each class in LVIS (since all classes may not have 10 examples). We use a batch size of $32$, learning rate of $52.5e-$ 5 for $46k$ iterations. We do not use Repeat Factor Sampling for fine-tuning. We sample $50$ categories for each training image, i.e $|S|=50$ for the FedLoss experiments. We derive negatives from pseudolabels with atleast $20\%$ confidence for the Psuedo-Negative experiment.